# A predictive modular approach to constraint satisfaction under uncertainty - with application to glycosylation in continuous monoclonal antibody biosimilar production

### v1

@@ 4.3 Constraint violation cost monitor
where $\rho_{monitor}\in(0,1)$ is a user-defined threshold, it indicates that the current constraint margin $\theta^{*}$ may no longer be optimal due changes in the uncertainties, and hence a new optimal $\theta$ should be learnt using the constraint adaptor.
@@ 5.1 Glycosolation in Continuous mAb Production
Glycosylation, a post-translational modification of proteins produced by mammalian cells, is a critical quality attribute (CQA) of monoclonal antibodies (mAbs). It has significant impact on properties such as the pharmacokinetics and pharmacodynamics of mAbs [53, 35]. The glycosylation attributes of biosimilar products are required to be within given reference ranges, set by regulatory agencies to ensure safety and efficacy [15, 33, 19]. If the reference ranges are exceeded, the biosimilar products need to be discarded or recycled, or it must be justified that the biosimilar has a similar effect as the reference product for approval. In any case, violating the given reference range will induce significant economic losses.
Controlling glycosylation such that it is maintained within a given reference range is a challenging task. This is in particular the case if the reactor is operated close to one of the reference limits, something which in general will be economically optimal provided violating the reference range can be avoided. Although many methods have been proposed for controlling glycosylation [72, 27, 32], most methods are based on open-loop control and hence sensitive to disturbances and model uncertainties. Recent advances in monitoring tools, including in-line Raman spectroscopy [55, 52, 18, 60, 44], and the development of mathematical models of glycosylation [58, 37, 73], have facilitated the application of closed-loop feedback control methods, such as MPC [74]. However, most control methods considered in bioprocessing do not explicitly consider constraint handling, neither do emerging explorative learning-based methods, and hence there is a need to add such a feature to the control system to avoid significant costs imposed by violating the imposed regulatory reference limits [69, 38, 29, 40].
We here formulate the glycosylation reference limits as soft constraints that can be temporarily violated but then with an economic cost that is penalized when determining the optimal constraint margin to operate at. The particular problem we consider is maintaining the level of the glycoform $G0$, reflecting the degree of glycosylation maturation [73, 53], within a given reference range in continuous production of mAbs. mAb-based biopharmaceuticals is becoming the dominating product in the pharmaceutical industry [67]. Continuous production of mAbs allows for higher productivity for an extended period of time as compared to more traditional batch processes [14].
@@ 5.2 Modeling
We consider glycosylation constraint satisfaction problem in continuous bioproduction, known as perfusion, as illustrated in Figure 3 [14, 9]. The mass balance equations [28] are given by
$\small\begin{split}\frac{d\text{V}[\text{M}_{\text{i}}]}{d\text{t}}&=\text{q}_{\text{M}_{\text{i}}}\text{V}[\text{X}_{\text{v}}]+\text{F}_{\text{{in}}}[\text{M}_{\text{{i,in}}}]-\text{F}_{\text{{out}}}[\text{M}_{\text{i}}],\\ \frac{d\text{V}[\text{mAb}]}{d\text{t}}&=\text{q}_{\text{{mAb}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{in}}}[\text{mAb}],\\ \frac{d\text{V}[\text{X}_{\text{v}}]}{d\text{t}}&=\text{q}_{\text{{biomass}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{bleed}}}[\text{X}_{\text{v}}],\end{split}$ (13)
with $\text{F}_{\text{{out}}}=\text{F}_{\text{{bleed}}}+\text{F}_{\text{{harvest}}}$. The unit and meaning of all variables and parameters are summarized in Table 1. To determine the production and consumption rates of the various external metabolites, we consider the metabolic network model developed in [25] based on experimental data presented in [24]. The model involves 23 extracellular metabolites and 126 reactions in mAb-producing mammalian CHO cells. Under a pseudo steady-state assumption [48], we have the external metabolite rates
@@
Having obtained the consumption rates from (13) and (14), we obtain the G0 level glycoform percentage $P_{G_{0}}$, i.e., the percentage of $G0$ and $G0F$ glycoforms among all fucosylated and afucosylated glycoforms, as a function of the consumption rate of Glucose $q_{Glc}$ using a static mathematical model presented in [72]. The static relationship is illustrated in Figure 4. Note that Glucose is the only fed carbon source in the considered example. The interested reader is referred to [72, 73] and [25] for details on the considered models.
@@ 5.3 Feedback control of glycosylation
As stated in Section 2, our goal is to minimize the economic cost induced by $P_{G_{0}}$ becoming lower than the imposed reference $G0$ glycoform percentage $P_{G_{0},opt}=35\%$. The most straightforward approach is to simply design the medium using available models and relationships, such that $P_{G_{0}}>P_{G_{0},opt}$, e.g., [72, 27, 32]. However, such an approach will be sensitive to model uncertainty and unmeasured disturbances. A simple extension is to combine the open-loop design with the proposed learning-based predictive filter to improve constraint satisfaction in the presence of uncertainty. However, due to the relatively slow dynamics of glycosylation, the collected data may not be sufficiently informative to learn the optimal constraint margins. We here propose a simple feedback control law combined with the learning predictive filter, and demonstrate its advantages compared with open-loop methods.
@@
where $u$ represents the percentage of low glucose solution; $u_{0}=85.7\%$ is the baseline input value corresponding to a predicted $P_{G_{0},opt}=35\%$ using the available model; $k=0.3$ is the proportional gain of the controller that adjusts $u$ based on measurements of $P_{G_{0}}$ with sampling time $T_{s}=5\text{ min}$. The sampling time is motivated by recent advances in monitoring tools such as in-line Raman spectroscopy [52] that can measure the glycosylation profile within minutes. The open-loop method, on the other hand, directly applies the medium consisting of $u_{0}=85.7\%$ of low glucose solution without adjusting the glucose concentration based on the measurements of $P_{G_{0}}$.
@@
Figure 6 illustrates the comparison of feedback control and open-loop method under uncertainty when regulating $P_{G_{0}}$ to $35\%$ from a medium with $u=100\%$ low glucose solution. It demonstrates that feedback control can reach the desired value $P_{G_{0},opt}$ within $50\text{ min}$, while the open-loop method requires more than $500\text{ min}$, i.e., about 10 times slower than that with feedback control. In the presence of uniformly distributed input uncertainty of 10% relative to the nominal control strategy (16), the feedback control demonstrates robustness to uncertainty, whereas the open-loop method exhibits large deviations.
Figure 7 demonstrates the ability of feedback control to compensate for uncertainty and to tightly control constraint violation costs when operating near the optimal operating point. In the presence of uniformly distributed input uncertainty of 10% relative to the nominal control strategy (16), the open-loop method leads to frequent and severe violations of the high-cost constraint $P_{G_{0}}>P_{G_{0},opt}$ due to the lack of feedback correction, while the feedback control strategy regulates $P_{G_{0}}$ around $P_{G_{0},opt}=35\%$ by utilizing in-line glycosylation measurements, resulting in significantly less severe high-cost constraint violations. We next combine the feedback control strategy with the proposed learning-based predictive filter to further mitigate the economic loss induced by high-cost constraint violations, under realistic settings.
@@ 5.4 Constraint handling using learning-based predictive filter
Consider now applying a learning-based controller that imitates the feedback control law (16) to control $G0$ level glycoform percentage $P_{G_{0}}$. To account for the explorative behavior of the learning-based controller, we assume the desired input is composed of the proportional control combined with an additive uniform stochastic part with a magnitude of 5% of the baseline input value $u_{0}$, that is,
@@
where $v_{u}\sim U(-0.05,0.05)$ and the other parameters are given in (16). We assume $P_{G_{0}}$ can be measured accurately by in-line measurements with sampling time $T_{s}=5\text{ min}$.
@@ 5.4.1 Uncertainties
    $\begin{split}\frac{d\text{V}[\text{M}_{\text{i}}]}{d\text{t}}&=\text{q}_{\text{M}_{\text{i}}}\text{V}[\text{X}_{\text{v}}]+\text{F}_{\text{{in}}}[\text{M}_{\text{{i,in}}}]-\text{F}_{\text{{out}}}[\text{M}_{\text{i}}]+d_{\text{M}_{\text{i}}},\\ \frac{d\text{V}[\text{mAb}]}{d\text{t}}&=\text{q}_{\text{{mAb}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{in}}}[\text{mAb}]+d_{[\text{mAb}]},\\ \frac{d\text{V}[\text{X}_{\text{v}}]}{d\text{t}}&=\text{q}_{\text{{biomass}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{bleed}}}[\text{X}_{\text{v}}]+d_{[\text{X}_{v}]};\end{split}$ (18)
@@ 5.4.2 Results
Due to the high experimental cost of obtaining a glycosylation model [73], we assume that a model to directly calculate $P_{G_{0}}$—as illustrated in Figure 4—is not available. Instead, we consider adapting the constraint margin of the glucose consumption rate $q_{Glc}$, whose behavior is well studied and which has a direct impact on $P_{G_{0}}$. Moreover, the empirical value corresponding to the desired $P_{G_{0}}$ can usually be obtained. We consider the following constraints in learning-based predictive filter:
@@
To achieve the objective, we apply the learning-based predictive filter described in Section 4. In the constraint adaptor, we introduce $\theta\in\mathbb{R}$ as the constraint margin of the high-cost constraints $g_{e}(x,u)$. The desired control signal $u^{d}\in\mathbb{R}$ is generated by the feedback controller (17). In the predictive filter, we apply a discrete-time model obtained by discretizing the nominal nonlinear model (13) with sampling time $T_{s}=5\text{ min}$. The prediction horizon is chosen to be $N=3$. The optimal constraint margin $\theta$ is learned online using the FDSA method, with step-size sequences that are adapted from [61] and [56]
@@
We next compare the performance of the feedback controller (17) with and without the learning-based predictive filter in the presence of the same uncertainties. As shown in Figure 10, the actual $G0$ level glycoform percentage $P_{G_{0}}$ is seldom below $P_{G_{0},opt}=35\%$ using the learning-based predictive filter with $\theta^{*}=0.026$, while the violation is severe and quite frequent using the feedback controller (17) only. Note that, due to the minimal modification to the desired input, the value of $q_{Glc}$ remains close to the desired $q_{Glc}$ when no filter is applied. Furthermore, the actual G0-level glycoform percentage, $P_{G_{0}}$, achieved using the learning-based predictive filter with the optimal constraint margin $\theta^{*}$, remains close to the target value $P_{G_{0},\text{opt}}=35\%$, exhibiting similar behavior to the case where only the proportional controller (17) is used. In this example, the desired inputs are frequently modified by the proposed method due to frequent constraint violations under uncertainty. In [68], we presented a fed-batch penicillin production example demonstrating that the desired control input is modified by the proposed method only when necessary. The proposed method ensures that performance is not significantly altered from the desired one while minimizing the impact of high-cost constraint violations under uncertainty.
@@
When applying the converged optimal constraint margin $\theta^{*}$, the violation cost should be monitored constantly such that the constraint margin can be adapted according to the change in the size and characteristics of uncertainties. For this purpose, we apply the proposed constraint violation cost monitor with the number of samples for averaging $n_{monitor}=48$ and threshold $\rho_{monitor}=10\%$. As illustrated in Figure 8 (b), within the first 10 monitoring counts, the monitored average constraint violation cost $J_{monitor}$ falls within the interval $(1-\rho_{monitor})J^{*},(1+\rho_{monitor})J^{*}]$ (red dashed lines) where $J^{*}=0.53$ is the average constraint violation cost corresponding to the converged $\theta^{*}=0.026$. This implies that the condition (12) is satisfied and the constraint margin $\theta^{*}$ is still optimal. After the $10^{\text{th}}$ monitoring count, we reduce the size of the disturbances described in Section 5.4.1 by $30\%$. As illustrated in Figure 8 (b), at the $11^{th}$ monitoring count, the monitored cost $J_{monitor}$ falls beyond the red dashed lines, meaning that the monitor detects that $|{J_{monitor}-J^{*}}|>\rho_{monitor}|{J^{*}}|$. This implies that the current constraint margin $\theta$ may no longer be optimal due to the change in the uncertainties. Therefore, a new optimal $\theta$ should be learned using the constraint adaptor. As shown in Figure 8 (c), a new optimal margin $\theta=0.021$ is learned, starting from the previous optimal margin $\theta^{*}=0.026$, and a lower long-term cost violation cost $0.38$ is reached. After converging to the new constraint margin, the violation cost is constantly monitored again by the violation cost monitor, as illustrated in Figure 8 (d). This shows that the proposed method has the ability to adjust the optimal constraint margin according to the detection of a change in the uncertainties.

### v3

@@ 4.3 Constraint violation cost monitor
where $\rho_{monitor}\in(0,1)$ is a user-defined threshold, it indicates that the current constraint margin $\theta^{*}$ may no longer be optimal due to changes in the uncertainties, and hence a new optimal $\theta$ should be learnt using the constraint adaptor.
@@ 4.4 Summary of learning-based predictive filter
4.5 Guideline for parameter selection
The parameters of the proposed method can be selected as follows. The initial constraint margin $\theta_{0}$ is typically set to $\theta_{0}=0$, after which the constraint adapter learns the optimal margin from online data. The prediction horizon $N$ and sampling time $T_{s}$ in the predictive filter are chosen based on the available sampling rate and computational capacity, with larger $N$ improving prediction accuracy but increasing computational load. The step-size sequences $\{a_{j}\}$ and $\{c_{j}\}$ follow standard FDSA guidelines [56, 57] and may be tuned heuristically [61] if needed. The convergence threshold $\rho_{\text{learning}}$ determines when the margin is considered converged, where smaller values lead to slower but more precise convergence. The averaging window $n_{\text{monitor}}$ should be sufficiently large to reduce stochastic variation in the violation cost, while the detection threshold $\rho_{\text{monitor}}$ specifies when the margin should be re-learned. Smaller $\rho_{\text{monitor}}$ values increase sensitivity to changes in the underlying uncertainty.
@@ 5.1 Glycosolation in Continuous mAb Production
Glycosylation, a post-translational modification of proteins produced by mammalian cells, is a critical quality attribute (CQA) of monoclonal antibodies (mAbs). It has significant impact on properties such as the pharmacokinetics and pharmacodynamics of mAbs [53, 35]. The glycosylation attributes of biosimilar products are required to be within given reference ranges - so-called specifications - to ensure product safety and efficacy [15, 33, 19]. If the reference ranges are exceeded, the biosimilar products needs to be rejected or recycled, or it must be justified that the biosimilar has a similar effect as the reference product for approval. In any case, violating the given reference range will induce significant economic losses.
Controlling glycosylation such that it is maintained within a given reference range is a challenging task. This is particularly the case in the production of biopharmaceuticals, where product quality can be affected by medium components or physical conditions when the reactor operates close to one of the reference limits—a situation that is generally economically optimal provided specification violations can be avoided. Although many methods have been proposed for controlling glycosylation [71, 27, 32], most methods are based on open-loop control and hence sensitive to disturbances and model uncertainties. Recent advances in monitoring tools, including in-line Raman spectroscopy [55, 52, 18, 60, 44], and the development of mathematical models of glycosylation [58, 37, 72], have facilitated the application of closed-loop feedback control methods, such as MPC [73]. However, most control methods considered in bioprocessing do not explicitly consider constraint handling, neither do emerging explorative learning-based methods, and hence there is a need to add such a feature to the control system to avoid significant costs imposed by violating the imposed regulatory reference limits [68, 38, 29, 40].
We here formulate the glycosylation reference limits as soft constraints that can be temporarily violated but then with an economic cost that is penalized when determining the optimal constraint margin to operate at. The particular problem we consider is maintaining the level of the glycoform $G0$, reflecting the degree of glycosylation maturation [72, 53], within a given reference range during continuous mAb production, which enables higher productivity over extended operating periods compared with traditional batch processes [14]. Note that product specifications on biopharmaceuticals typically are defined based on a batch averaged quality, but that the introduction of continuous perfusion processes makes it feasible to impose quality specifications on the instantaneous production. In this case study we consider the latter case. We assume that in-line measurements of glucose consumption rates are available and that the level of glycosylation can be estimated based on these measurements and the model presented in [71, 72].
@@ 5.2 Modeling
We consider glycosylation constraint satisfaction problem in continuous bioproduction, known as perfusion operation, as illustrated in Figure 3 [14, 9]. The mass balance equations [28] are given by
$\small\begin{split}\frac{d\text{V}[\text{M}_{\text{i}}]}{d\text{t}}&=\text{q}_{\text{M}_{\text{i}}}\text{V}[\text{X}_{\text{v}}]+\text{F}_{\text{{in}}}[\text{M}_{\text{{i,in}}}]-\text{F}_{\text{{out}}}[\text{M}_{\text{i}}],\\ \frac{d\text{V}[\text{mAb}]}{d\text{t}}&=\text{q}_{\text{{mAb}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{out}}}[\text{mAb}],\\ \frac{d\text{V}[\text{X}_{\text{v}}]}{d\text{t}}&=\text{q}_{\text{{biomass}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{bleed}}}[\text{X}_{\text{v}}],\end{split}$ (13)
with $\text{F}_{\text{{out}}}=\text{F}_{\text{{bleed}}}+\text{F}_{\text{{harvest}}}$. The unit and meaning of all variables and parameters are summarized in Table 1. To determine the production and consumption rates of the various external metabolites, we consider the metabolic network model developed in [25] based on experimental data presented in [24]. The model involves 23 extracellular metabolites and 126 reactions in mAb-producing mammalian CHO cells. Under a pseudo steady-state assumption [48], the external metabolite rates are given by
@@
Having obtained the consumption rates from (13) and (14), the steady-state G0-level glycoform percentage $P_{G0}$—representing the proportion of G0 and G0F among all fucosylated and afucosylated glycans—is estimated as a function of the glucose consumption rate $q_{\mathrm{Glc}}$ using a steady-state model from Zhang et al. [71], assuming glucose is the only fed carbon source:
$P_{G0}(q_{\mathrm{Glc}})=\left[1-\frac{P_{G1F}(q_{\mathrm{Glc}})+P_{G2F}(q_{\mathrm{Glc}})+P_{G1}(q_{\mathrm{Glc}})}{P_{\mathrm{nonHM}}(q_{\mathrm{Glc}})}\right]\times 100\%,$ (16)
where the galactosylated glycoform fractions $P_{G1F}$, $P_{G2F}$, and $P_{G1}$ follow Michaelis–Menten-type functions of $q_{\mathrm{Glc}}$. The non–high-mannose fraction is given by
$P_{\mathrm{nonHM}}(q_{\mathrm{Glc}})=P_{\max}\,\Phi_{\mathrm{nac}}(q_{\mathrm{Glc}})\,\Psi_{\mathrm{man}}(q_{\mathrm{Glc}}),$ (17)
with $\Phi_{\mathrm{nac}}(\cdot)$ and $\Psi_{\mathrm{man}}(\cdot)$ representing saturation and inhibition effects associated with N-acetylglucosamine (GlcNAc) and mannose precursors respectively, each described by Michaelis–Menten-type kinetics. $P_{\max}$ represents the upper limit of non-high-mannose glycans. Parameter values are adopted from Zhang et al. [72]. The static relationship is illustrated in Figure 4. It serves as a static proxy that relates the steady-state glycosylation quality to glucose consumption. The interested reader is referred to [71, 72] and [25] for details on the considered model.
In the predictive filter, we assume that (i) in-line measurements of consumption rates are available at a 5-min sampling interval (e.g., via Raman spectroscopy), and (ii) the steady-state glycosylation pattern can be estimated from these measurements using the static model in Zhang et al. [71, 72]. This static model has been experimentally validated in continuous mAb production [71], making such measurements realistic and implementable in practice. If the glycosylation model is inaccurate, data from previous experiments could be used to construct confidence intervals or quantify estimation uncertainty, enabling a robust extension of the proposed method. Addressing such extensions, however, lies beyond the scope of the present study.
@@ 5.3 Feedback control of glycosylation
As stated in Section 2, our goal is to minimize the economic cost induced by $P_{G_{0}}$ becoming lower than the imposed reference $G0$ glycoform percentage $P_{G_{0},opt}=35\%$. The most straightforward approach is to simply design the medium using available models and relationships, such that $P_{G_{0}}>P_{G_{0},opt}$, e.g., [71, 27, 32]. However, such an approach will be sensitive to model uncertainty and unmeasured disturbances. A straightforward extension is to combine the open-loop design with the proposed learning-based predictive filter to enhance constraint satisfaction under uncertainty. However, because metabolic dynamics are typically slow, the collected data may not be sufficiently informative for learning the optimal constraint margins. To address this, we propose a simple feedback control law integrated with the learning-based predictive filter, demonstrating improved performance under uncertainty compared with open-loop approaches.
@@
where $u$ represents the percentage of low glucose solution; $u_{0}=85.7\%$ is the baseline input value corresponding to a predicted $P_{G_{0},opt}=35\%$ using the available model; $k=0.3$ is the proportional gain of the controller that adjusts $u$ based on estimates of $P_{G_{0}}$ with sampling time $T_{s}=5\text{ min}$. The sampling time is motivated by recent advances in monitoring tools such as in-line Raman spectroscopy [52] that allows estimation of the glycosylation profile at steady state based on in-line measurements of the glucose consumption rate within minutes. The open-loop method, on the other hand, directly applies the medium consisting of $u_{0}=85.7\%$ of low glucose solution without adjusting the glucose concentration based on the estimates of $P_{G_{0}}$. Note that the controller (18) effectively is a model based feedback controller from in-line measurements of the glucose consumption rate.
@@
We simulate the system under uniformly distributed input uncertainty of 10% relative to the nominal control strategy (18). Figure 6 compares the feedback and open-loop methods when regulating the steady-state $P_{G_{0}}$ estimate to $P_{G_{0},opt}=35\%$ from a medium with $u=100\%$ low-glucose solution. The feedback controller reaches the target $P_{G_{0},opt}$ within $50$ min, whereas the open-loop approach requires more than $500$ min. Figure 7 further shows that, under the same 10% input uncertainty, the feedback control maintains robustness and limits constraint violations, while the open-loop method exhibits large deviations and frequent high-cost violations of the constraint $P_{G_{0}}>P_{G_{0},opt}$. The results highlight the importance of feedback control for regulating glycosylation and ensuring consistent product quality under uncertainty.
We next combine the feedback control strategy with the proposed learning-based predictive filter to further mitigate the economic loss induced by high-cost constraint violations, under realistic settings.
@@ 5.4 Constraint handling using learning-based predictive filter
Consider now applying a learning-based controller that imitates the feedback control law (18) to control steady-state $G0$ level glycoform percentage estimate $P_{G_{0}}$. To account for the explorative behavior of the learning-based controller, we assume the desired input is composed of the proportional control combined with an additive uniform stochastic part with a magnitude of 5% of the baseline input value $u_{0}$, that is,
@@
where $v_{u}\sim U(-0.05,0.05)$ and the other parameters are given in (18). We assume the $P_{G_{0}}$ can be predicted accurately by in-line measurements with sampling time $T_{s}=5\text{ min}$.
@@ 5.4.1 Uncertainties
    $\begin{split}\frac{d\text{V}[\text{M}_{\text{i}}]}{d\text{t}}&=\text{q}_{\text{M}_{\text{i}}}\text{V}[\text{X}_{\text{v}}]+\text{F}_{\text{{in}}}[\text{M}_{\text{{i,in}}}]-\text{F}_{\text{{out}}}[\text{M}_{\text{i}}]+d_{\text{M}_{\text{i}}},\\ \frac{d\text{V}[\text{mAb}]}{d\text{t}}&=\text{q}_{\text{{mAb}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{in}}}[\text{mAb}]+d_{[\text{mAb}]},\\ \frac{d\text{V}[\text{X}_{\text{v}}]}{d\text{t}}&=\text{q}_{\text{{biomass}}}\text{V}[\text{X}_{\text{v}}]-\text{F}_{\text{{bleed}}}[\text{X}_{\text{v}}]+d_{[\text{X}_{v}]},\end{split}$ (20)
@@ 5.4.2 Results
In the predictive filter we impose a constraint on the Glucose consumption rate $q_{\mathrm{Glc}}$, which is the measured variable and, furthermore, directly correlated with $P_{G_{0}}$. This choice is partly motivated by the fact that the dynamics of the glucose consumption rate are well studied, making a predictive model for $q_{\mathrm{Glc}}$ readily available, whereas obtaining a dynamic model for $P_{G_{0}}$ would require extensive and costly experiments. In the violation cost monitor, the constraint violation cost is evaluated based on a steady-state estimate of $P_{G_{0}}$ using the glucose consumption rate measurements combined with the static model in Zhang et al. [71, 72].
We hence consider the following constraints in the learning-based predictive filter:
@@
To achieve the objective, we apply the learning-based predictive filter described in Section 4. In the constraint adaptor, we introduce $\theta\in\mathbb{R}$ as the constraint margin of the high-cost constraints $g_{e}(x,u)$. The desired control signal $u^{d}\in\mathbb{R}$ is generated by the feedback controller (19). In the predictive filter, we apply a discrete-time model obtained by discretizing the nominal nonlinear model (13) with sampling time $T_{s}=5\text{ min}$. The prediction horizon is chosen to be $N=3$. The optimal constraint margin $\theta$ is learned online using the FDSA method, with step-size sequences that are adapted from [61] and [56] [^3] We follow the functional form and parameter choices of [61] with minor rescaling to achieve a practical convergence rate in the simulation setup.
@@
We next compare the performance of the feedback controller (19) with and without the learning-based predictive filter in the presence of the same uncertainties. As shown in Figure 10, the $G0$ level glycoform percentage $P_{G_{0}}$ is seldom below $P_{G_{0},opt}=35\%$ using the learning-based predictive filter with $\theta^{*}=0.026$, while the violation is severe and quite frequent using the feedback controller (19) only. Note that, due to the minimal modification to the desired input, the value of $q_{Glc}$ remains close to the desired $q_{Glc}$ when no filter is applied. Furthermore, the G0-level glycoform percentage, $P_{G_{0}}$, achieved using the learning-based predictive filter with the optimal constraint margin $\theta^{*}$, remains close to the target value $P_{G_{0},\text{opt}}=35\%$, exhibiting similar behavior to the case where only the proportional controller (19) is used. In this example, the desired inputs are frequently modified by the proposed method due to frequent constraint violations under uncertainty. In [67], we presented a fed-batch penicillin production example demonstrating that the desired control input is modified by the proposed method only when necessary. The proposed method ensures that performance is not significantly altered from the desired one while minimizing the impact of high-cost constraint violations under uncertainty.
@@
When applying the converged optimal constraint margin $\theta^{*}$, the violation cost should be monitored constantly such that the constraint margin can be adapted according to the change in the size and characteristics of uncertainties. For this purpose, we apply the proposed constraint violation cost monitor with the number of samples for averaging $n_{monitor}=48$ and threshold $\rho_{monitor}=10\%$. As illustrated in Figure 8 (b), within the first 10 monitoring counts, the monitored average constraint violation cost $J_{monitor}$ falls within the interval $(1-\rho_{monitor})J^{*},(1+\rho_{monitor})J^{*}]$ (red dashed lines) where $J^{*}=0.53$ is the average constraint violation cost corresponding to the converged $\theta^{*}=0.026$. This implies that the condition (12) is satisfied and the constraint margin $\theta^{*}$ is still optimal. After the $10^{\text{th}}$ monitoring count, we reduce the size of the disturbances described in Section 5.4.1 by $30\%$. As illustrated in Figure 8 (b), at the $11^{th}$ monitoring count, the monitored cost $J_{monitor}$ falls beyond the red dashed lines, meaning that the monitor detects that $|{J_{monitor}-J^{*}}|>\rho_{monitor}|{J^{*}}|$. This implies that the current constraint margin $\theta$ may no longer be optimal due to the change in the uncertainties. Therefore, a new optimal $\theta$ should be learned using the constraint adaptor. As shown in Figure 8 (c), a new optimal margin $\theta=0.021$ is learned, starting from the previous optimal margin $\theta^{*}=0.026$, and a lower long-term cost violation cost $0.38$ is reached. After converging to the new constraint margin, the violation cost is constantly monitored again by the violation cost monitor, as illustrated in Figure 8 (d). This shows that the proposed method has the ability to adjust the optimal constraint margin in response to the detection of a change in the uncertainties.